OpenStudy (anonymous):

X + y = 4 X - 4 = y/2

OpenStudy (anonymous):

what r u trying to solve?

OpenStudy (anonymous):

Solving for x and y?

OpenStudy (radar):

Take that 2nd equation and multiply thru by 2, getting 2x -8 = y now rearrange it (using standard rules of algebra) 2x - y = 8 Now look at what you now have: x + y = 4 2x - y = 8 Now add those two and solve for x.